What volume is occupied by 6.00 g of oxygen gas (O2) under a pressure of 3.60 atm and a temperature of 293 K ? Use 1atm=1.01×105Pa if needed.
Answer #1

Here,

let the volume is V

Using ideal gas equation

P V = n * R * T

3.6 *1.01 *10^5 * V = 6/32 * 8.314 * 293

solving for V

V = 1.26 * 10^-3 m^3

the volume of the 6 g is 1.26 * 10-3 m^3